# Rate, Explain and Cite (REC): Enhanced Explanation and Attribution in Automatic Evaluation by Large Language Models

### v2

@@ B.4 Inline Citation
{  "citations": [    {      "context_id": "1233",         "claim": "Photosynthesis is a process         that converts carbon dioxide         and water into glucose and         oxygen using sunlight and         chlorophyll."    }  ]}
@@ B.5 Inline Citation with Context Snippet
{  "citations": [    {      "context_id": "1233",       "claim": "Photosynthesis is a process         that converts carbon dioxide         and water into glucose and         oxygen using sunlight and         chlorophyll.",       "snippet": "Photosynthesis is the         process by which green         plants and some other         organisms utilize sunlight         to synthesize their food. This         remarkable process involves         the conversion of carbon         dioxide and water into         glucose and oxygen,         facilitated by chlorophyll"    }  ]}
@@ Appendix C REC-Data Details
For general instruction tasks, we include summarization on ABCD (Chen et al., 2021), a dataset containing customer support dialogue transcripts, in five languages. For pointwise evaluation, we leverage HelpSteer2 (Wang et al., 2024c) and perform synthetic data generation (Section 4.2) on the (dialogue, summary) pairs from ABCD, and QA pairs from FloDial. For pairwise and open-ended evaluation tasks, we use Auto-j (Li et al., 2023a), Shepherd (Wang et al., 2023b), Skywork  [^10] https://huggingface.co/datasets/Skywork/Skywork-Reward-Preference-80K-v0.1, HelpSteer2111111https://huggingface.co/datasets/nvidia/HelpSteer2, OffsetBias [^12] https://huggingface.co/datasets/NCSOFT/offsetbias, and Code Preference [^13] https://huggingface.co/datasets/Vezora/Code-Preference-Pairs. For citations, we perform synthetic data generation (Section 4.2) on FloDial for both RAG citations and content quality citations, and on ABCD, Auto-j, Shepherd, and Email Thread Summary  [^14] https://www.kaggle.com/datasets/marawanxmamdouh/email-thread-summary-dataset/data for content quality citations.
@@ Appendix D Synthetic Data Generation Prompt Example
D.1 Pointwise Evaluation ("Faithfulness metric")
You will be given a source text given to anotherLLM, and that LLM’s answer.Your task is to rate that LLM’s answeron one quality metric.Please make sure you read and understandthese instructions carefully. Pleasekeep this document open while reviewing,and refer to it as needed.Evaluation criteria:metric_name = ’Faithfulness’metric_scale = ’Yes/No’metric_description = ""The generated answer only contains truthfulcontent, and does not contain inventedor misleading facts that arenot supported by the context.""Evaluation steps:1. Read the source text.2. Read the other LLM’s answer and compareit to the source text. Check if theevaluation criteria is met.3. If the evaluation criteria is met,answer Yes; if the evaluation criteriais not met, answer No.4. Create a JSON response{    "metriclabel": <Yes|No only>,    "justification":    <explanation for the score>}Example:Source text given to another LLM:{query_with_context}The other LLM’s answer:{answer}### Response(JSON Only):
@@ Appendix E Remark on Why The Three Metrics in ALCE are All Important
Regarding why the three criteria are equally important, we encourage the reader to consider the user’s perspective. An LLM’s citation response cannot be deemed good if it is lacking in any of the three criteria: fluency, answer correctness, and citation accuracy. To illustrate this, we aim to provide an example of a good citation response that scores well on all three criteria, as well as three examples that each fail in one of the criteria (assuming the same RAG QA setting as in Appendix B).
@@ Appendix F ABCD Content Quality Citation Human Evaluation Instruction
Annotation instruction: You will be evaluating attributed judgments made by an LLM autoevaluator for answers generated by another LLM against three metrics: instruction-following, completeness, and faithfulness. Your task is to evaluate LLM autoevaluator’s answer by completing the following THREE tasks: *you can skip if the output shows “invalid json”.
  • Metric answer correctness: Please label “yes”/“no” by evaluating whether the metric “answer” in autoevaluator’s output is correct or not, based on the provided conversation, task prompt, and summary.
  • Metric explanation correctness: Please label “yes”/“no’ by evaluating whether the “feedback” in autoevaluator’s output is correct or not, based on the provided conversation, task prompt, and summary. Only proceed to citation annotation (next task) if you label “yes” in this task, namely the explanation is correct.
  • Citation annotation: In “statements”, you will find different “statement_string” extracted from the “feedback”. Please provide citation for each “statement_string” by directly copying relevant sentences (Please cite full sentences, not substrings) from “task prompt” and “chat” columns (Please DON’T cite from “summary”). If the “statement_string” is hallucinated and doesn’t make sense in the first place, please put “[<halu>]” as the citation to signal.
@@ Appendix G Prompt for LLM-AggreFact Eval
Determine whether the provided claim is consistentwith the corresponding document. Consistency inthis context implies that all information presentedin the claim is substantiated by the document.If not, it should be considered inconsistent.Document: {doc}Claim: {claim}Please assess the claim’s consistency with thedocument by responding with either "yes" or "no".Answer (yes|no only):

### v3

@@ B.4 Inline Citation
  "citations": [
      "context_id": "1233",
         "claim": "Photosynthesis is a
         process that converts carbon
         dioxide and water into glucose
         and oxygen using sunlight and
         chlorophyll."
@@ B.5 Inline Citation with Context Snippet
  "citations": [
      "context_id": "1233",
       "claim": "Photosynthesis is a
         process that converts carbon
         dioxide and water into glucose
         and oxygen using sunlight and
         chlorophyll.",
       "snippet": "Photosynthesis is the
         process by which green
         plants and some other
         organisms utilize sunlight
         to synthesize their food. This
         remarkable process involves
         the conversion of carbon
         dioxide and water into
         glucose and oxygen,
         facilitated by chlorophyll"
@@ Appendix C REC-Data Details
For general instruction tasks, we include summarization on ABCD (Chen et al., 2021), a dataset containing customer support dialogue transcripts, in five languages. For pointwise evaluation, we leverage HelpSteer2 (Wang et al., 2024c) and perform synthetic data generation (Section 4.2) on the (dialogue, summary) pairs from ABCD, and QA pairs from FloDial. For pairwise and open-ended evaluation tasks, we use Auto-j (Li et al., 2023a), Shepherd (Wang et al., 2023b), Skywork  [^7] https://huggingface.co/datasets/Skywork/Skywork-Reward-Preference-80K-v0.1, HelpSteer2888https://huggingface.co/datasets/nvidia/HelpSteer2, OffsetBias [^9] https://huggingface.co/datasets/NCSOFT/offsetbias, and Code Preference [^10] https://huggingface.co/datasets/Vezora/Code-Preference-Pairs. For citations, we perform synthetic data generation (Section 4.2) on FloDial for both RAG citations and content quality citations, and on ABCD, Auto-j, Shepherd, and Email Thread Summary  [^11] https://www.kaggle.com/datasets/marawanxmamdouh/email-thread-summary-dataset/data for content quality citations.
@@ Appendix D Synthetic Data Generation Prompt Example
D.1 Pointwise Evaluation (”Faithfulness metric”)
You will be given a source text given to
another LLM, and that LLM’s answer.
Your task is to rate that LLM’s answer
on one quality metric.
Please make sure you read and understand
these instructions carefully. Please
keep this document open while reviewing,
and refer to it as needed.
Evaluation criteria:
metric_name = ’Faithfulness’
metric_scale = ’Yes/No’
metric_description = ""
The generated answer only contains truthful
content, and does not contain invented
or misleading facts that are
not supported by the context.
Evaluation steps:
1. Read the source text.
2. Read the other LLM’s answer and compare
it to the source text. Check if the
evaluation criteria is met.
3. If the evaluation criteria is met,
answer Yes; if the evaluation criteria
is not met, answer No.
4. Create a JSON response
    "metriclabel": <Yes|No only>,
    "justification":
    <explanation for the score>
Example:
Source text given to another LLM:
{query_with_context}
The other LLM’s answer:
{answer}
### Response(JSON Only):
@@ Appendix E Remark on Why The Three Metrics in ALCE are All Important
Regarding why the three criteria are equally important, we encourage the reader to consider from an user’s perspective. An LLM’s citation response cannot be deemed good if it is lacking in any of the three criteria: fluency, answer correctness, and citation accuracy. To illustrate this, we aim to provide an example of a good citation response that scores well on all three criteria, as well as three examples that each fail in one of the criteria (assuming the same RAG QA setting as in Appendix B).
@@ Appendix F ABCD Content Quality Citation Human Evaluation Instruction
Annotation instruction: You will be evaluating attributed judgments made by an LLM auto-evaluator for answers generated by another LLM against three metrics: instruction-following, completeness, and faithfulness. Your task is to evaluate LLM auto-evaluator’s answer by completing the following THREE tasks: *you can skip if the output shows “invalid json”.
  • Metric answer correctness: Please label “yes”/“no” by evaluating whether the metric “answer” in auto-evaluator’s output is correct or not, based on the provided conversation, task prompt, and summary.
  • Metric explanation correctness: Please label “yes”/“no’ by evaluating whether the “feedback” in auto-evaluator’s output is correct or not, based on the provided conversation, task prompt, and summary. Only proceed to citation annotation (next task) if you label “yes” in this task, namely the explanation is correct.
  • Citation annotation: In “statements”, you will find different “statement_string” extracted from the “feedback”. Please provide citation for each “statement_string” by directly copying relevant sentences (Please cite full sentences, not substrings) from “task prompt” and “chat” columns (Please DON’T cite from “summary”). If the “statement_string” is hallucinated and doesn’t make sense in the first place, please put “[¡halu¿]” as the citation to signal.
@@ Appendix G Prompt for LLM-AggreFact Eval
Determine whether the provided claim is
consistent with the corresponding document.
Consistency in this context implies that
all information presented in the claim is
substantiated by the document. If not,
it should be considered inconsistent.
Document: {doc}
Claim: {claim}
Please assess the claim’s consistency
with the document by responding with
either "yes" or "no".
Answer (yes|no only):